OpenStudy (anonymous):

Determine the steady-state transfer function of the following transform transfer function? G(s)=5s/(s+1)^(2)

OpenStudy (anonymous):

I would like actually to know a bit about these myself so maybe we can help each other. If I understand it correctly, then Y(s)/X8s) = G(s) -> (s+1)^2 Y(s) = 5s X(s) -> (s^2+2s+1) Y(s) = 5s (X(s) so y''(t) + 2y'(t) + y(t) = 5x'(t) Is that right?

OpenStudy (anonymous):

I am printing it to read....tks

OpenStudy (anonymous):

OK, let me study them... This explains the other post u made as well, yes?

OpenStudy (anonymous):

That's all I have.Have fun.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Ok, I got them all, 1 and 2 are the same, right?

OpenStudy (anonymous):

I am missing page 263....u have?

OpenStudy (anonymous):

Ta!